# HypeRL: Parameter-Informed Reinforcement Learning for Parametric PDEs

### v1

@@ 2.2.1 Twin-Delayed Deep Deterministic Policy Gradient
$\begin{split}Q(\bm{y}_{k},\bm{u}_{k})&=\mathbb{E}_{\pi}[\Sigma_{j=0}^{H}\gamma^{j}r_{k+j}|\text{Y}_{k}=\bm{y}_{k},\text{U}_{k}=\bm{u}_{k}]\\ &=\mathbb{E}_{\pi}[r_{k}+\Sigma_{j=1}^{H}\gamma^{j}r_{k+j}|\text{Y}_{k}=\bm{y}_{k},\text{U}_{k}=\bm{u}_{k}]\\ &=\mathbb{E}_{\pi}[r_{k}+\gamma\max_{\bm{u}_{k}\in U_{\text{ad}}}Q(\text{Y}_{k+1},\cdot)|\text{Y}_{k}=\bm{y}_{k},\text{U}_{k}=\bm{u}_{k}]\,,\\ \end{split}$ (14)
@@
$Q(\bm{y}_{k},\bm{u}_{k})\leftarrow Q(\bm{y}_{k},\bm{u}_{k})+\alpha\Big{(}r_{k}+\gamma\max_{\bm{u}_{k}\in U_{\text{ad}}}Q(\bm{y}_{k+1},\cdot)-Q(\bm{y}_{k},\bm{u}_{k})\Big{)}\,,$ (15)
@@
$Q(\bm{y}_{k},\bm{u}_{k})\leftarrow Q(\bm{y}_{k},\bm{u}_{k})+\alpha\Big{(}r_{k}+\gamma Q(\bm{y}_{k+1},\bm{u}_{k+1})-Q(\bm{y}_{k},\bm{u}_{k})\Big{)}\,,$ (16)
@@
TD3 relies on a memory buffer $\mathcal{D}$ to store the interaction data $(\bm{y}_{k},\bm{u}_{k},r_{k},\bm{y}_{k+1})$ for all time-steps $k$. Given a randomly-sampled batch of interaction tuples, we can employ Equation (16) as a loss function for updating the parameters $\bm{\theta}_{Q}$ of the action-value function $Q(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q})$ as:
$\begin{split}\mathcal{L}(\bm{\theta}_{Q})&=\mathbb{E}_{\bm{y}_{k},\bm{u}_{k},\bm{y}_{k+1},r_{k}\sim\mathcal{D}}[(r_{k}+\gamma\bar{Q}(\bm{y}_{k+1},\bm{u}_{k+1};\bm{\theta}_{\bar{Q}})-Q(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q}))^{2}]\\ &=\mathbb{E}_{\bm{y}_{k},\bm{u}_{k},\bm{y}_{k+1},r_{k}\sim\mathcal{D}}[(\underbrace{r_{k}+\gamma\bar{Q}(\bm{y}_{k+1},\bar{\pi}(\bm{y}_{k+1};\bm{\theta}_{\bar{\pi}})+\bm{\epsilon};\bm{\theta}_{\bar{Q}})}_{\text{target value}}-Q(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q}))^{2}]\,,\end{split}$ (17)
where the so-called target networks $\bar{Q}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{\bar{Q}})$ and $\bar{\pi}(\bm{y}_{k};\bm{\theta}_{\bar{\pi}})$ are copies of $Q(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q})$ and $\pi(\bm{y}_{k};\bm{\theta}_{\pi})$, respectively, with frozen parameters, i.e., they are not updated in the backpropagation step to improve the stability of the training. We indicate with $\bm{\epsilon}\sim\text{clip}(\mathcal{N}(\bm{0},\bar{\bm{\sigma}}),-c,c)$ the noise added to estimate the action value in the interval $[-c,c]$ around the target action. To reduce the problem of overestimation of the target Q-values, TD3 estimates two independent action-value functions, namely $Q_{1}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q_{1}})$ and $Q_{2}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q_{2}})$, and two target action-value functions $\bar{Q}_{1}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{\bar{Q}_{1}})$ and $\bar{Q}_{2}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{\bar{Q}_{2}})$, and computes the target value for regression (see (17)) as:
$\underbrace{r_{k}+\gamma\min_{i=1,2}\bar{Q}_{i}(\bm{y}_{k+1},\bm{u}_{k+1};\bm{\theta}_{\bar{Q}_{i}}).}_{\text{target value}}$ (18)
@@
$\mathcal{L}(\bm{\theta}_{\pi})=\mathbb{E}_{\bm{y}_{k}\sim\mathcal{D}}[-\nabla_{\bm{u}_{k}}Q_{1}(\bm{y}_{k},\pi(\bm{y}_{k};\bm{\theta}_{\pi});\bm{\theta}_{Q_{1}})].$ (19)
@@
$\begin{split}\bm{\theta}_{\bar{Q}_{1}}&=\rho\bm{\theta}_{Q_{1}}+(1-\rho)\bm{\theta}_{\bar{Q}_{1}}\,,\\ \bm{\theta}_{\bar{Q}_{2}}&=\rho\bm{\theta}_{Q_{2}}+(1-\rho)\bm{\theta}_{\bar{Q}_{1}}\,,\\ \bm{\theta}_{\bar{\pi}}&=\rho\bm{\theta}_{\pi}+(1-\rho)\bm{\theta}_{\bar{\pi}}\,,\\ \end{split}$ (20)
@@ 2.3 Hypernetworks
A hypernetwork [36] is a neural network (NN) that generates the weights and biases of another NN, often referred to as main or primary network. Formally, a hypernetwork $h:Z\subset\mathbb{R}^{|\bm{z}|}\to\mathbb{R}^{|\bm{\theta}_{f}|}$ learns the parameters $\bm{\theta}_{f}$ of the main network $f:X\subset\mathbb{R}^{|\bm{x}|}\to W\subset\mathbb{R}^{|\bm{w}|}$. If we consider a standard supervised learning regression task, and we assume the availability of a dataset of $N$ input-output pairs $[(\bm{x}^{(1)},\bm{w}^{(1)}),\cdots,(\bm{x}^{(N)},\bm{w}^{(N)})]$ and hypernetwork inputs $[\bm{z}^{(1)},\ldots,\bm{z}^{(N)}]$, we can write:
$\begin{split}\bm{\theta}_{f}&=h(\bm{z}^{(i)};\bm{\theta}_{h})\,,\\ \hat{\bm{w}}^{(i)}&=f(\bm{x}^{(i)};\bm{\theta}_{f})\,,\end{split}$ (21)
@@ 3 Methodology for HypeRL
Given all the elements introduced in the previous section, we are now ready to set our proposed strategy to address OC of parametric PDEs through RL.
@@ 3.1 Problem Settings
$\begin{split}\bm{y}_{k+1}&=\bm{T}(\bm{y}_{k},\bm{u}_{k};\bm{\mu})\,,\\ r_{k}&=R(\bm{y}_{k},\bm{u}_{k})\,,\\ \end{split}$ (23)
where the transition function is assumed to be deterministic and dependent on the PDE parameters $\bm{\mu}$. The transition function corresponds to the fully-discretized form of the state equation (see Equation (1)) obtained by introducing a suitable time-integration scheme over a partition of $[t_{0},t_{f}]$ made by $N_{t}$ time-steps $\{t_{k}\}_{k=0}^{N_{t}}$ such that the step size is $\Delta t=(t_{f}-t_{0})/N_{t}$. For example, by using an explicit Runge-Kutta scheme, we obtain:
$\begin{split}\bm{y}_{k+1}=\underbrace{\bm{y}_{k}+\Delta t\Phi(t_{k},\bm{y}_{k},\bm{u}_{k};\Delta t,\bm{F},\bm{\mu})}_{:=\,T(\bm{y}_{k},\bm{u}_{k};\bm{\mu})},\qquad k=0,\ldots,N_{t}-1,\end{split}$ (24)
where $\bm{y}_{k}\approx\bm{y}(t_{k})$ and $\bm{u}_{k}\approx\bm{u}(t_{k})$, and $\Phi$ denotes the integration method’s increment function related to the state equation. The reward function corresponds to the fully-discretized counterpart of the running cost in Equation (3) with opposite sign. Starting from (3), in the following, our goal is to steer the PDE state to a reference state while tracking a reference control signal. Therefore, we utilize the following reward function:
$R(\bm{y}_{k},\bm{u}_{k})=-\frac{1}{2}||\bm{y}_{k}-\bm{y}_{\text{ref}}||_{2}^{2}-\frac{\alpha}{2}||\bm{u}_{k}-\bm{u}_{\text{ref}}||_{2}^{2}=-\frac{1}{2}c_{1}-\frac{\alpha}{2}c_{2}\,,$ (25)
where $\bm{y}_{\text{ref}}$ and $\bm{u}_{\text{ref}}$ indicate the reference values for the state and the control input, and $\alpha$ is a scalar positive coefficient balancing the contribution of the two terms.
With reference to Algorithm 1, the agent-environment interaction scheme is organized in two loops: the outer loop indicates the training episode; instead, the inner loop refers to the number of time-steps for each episode. At the beginning of each new episode we sample a random initial condition and a PDE parameters value $\bm{\mu}$ to obtain the initial state. The agent utilizes the state at the current time-step and the PDE parameters to select a control input. The control input is fed to the transition and reward models to obtain the state at the next time-step and the reward. The tuples $(\bm{y}_{k},\bm{u}_{k},r_{k},\bm{y}_{k+1},\bm{\mu})$ collected at each time-step $k$ of the interaction are used to train the policy (and value function) deep NN (DNN) to maximize the expected cumulative reward at each episode.
Algorithm 1 Episodic RL for control of parametric PDEs
Initialize policy and value function DNN parameters $\bm{\theta}_{\pi}$ and $\bm{\theta}_{Q}$
@@
     Sample an initial condition and PDE parameter $\bm{\mu}$
@@
         Sample action from a policy $\bm{u}_{k}\sim\pi(\bm{y}_{k},\bm{\mu};\bm{\theta}_{\pi})+\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{\sigma})$
         Observe reward $r_{k}=R(\bm{y}_{k},\bm{u}_{k})$ and new state $\bm{y}_{k+1}=T(\bm{y}_{k},\bm{u}_{k};\bm{\mu})$
@@
              Update policy and value function using the tuple $(\bm{y}_{k},\bm{u}_{k},r_{k},\bm{y}_{k+1},\bm{\mu})$
@@
Differently from the majority of the literature on RL for OC of PDEs, we focus on devising a control strategy that is adaptable to changes of the systems’ dynamics deriving from variations of known parameters $\bm{\mu}$. We assume the agent to be able to observe the PDE state and the parameters $\bm{\mu}$. In this setting, we present a novel RL framework that can efficiently learn control policies for parametric PDEs from limited samples and that can generalize to new, unseen instances of the PDE parameters $\bm{\mu}$ (see Figure 1). In contrast with the widely-used concatenation of information in the agent’s state, to enhance the sample-efficiency and generalization capabilities of RL agents, we propose a parameter-informed HypeRL architecture relying on hypernetworks (see Section 2.3). Hypernetworks allow us to express the weights and biases of the value function and policy as functions of the PDE parameters $\bm{\mu}$ (see Figure 2). This new paradigm for encoding the information of the PDE parameters drastically improves the performance of the RL agent in terms of total cumulative reward, sample efficiency, and generalization with respect to traditional RL approaches.
3.2 HypeRL TD3
@@
In this work, we enhance the TD3 algorithm (see Section 2.2.1) with hypernetworks. However, our method can be easily and directly applied to other RL algorithms, such as PPO and SAC. Analogously to the TD3 algorithm, we rely on the estimation of two action-value functions $Q_{1}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q_{1}})$ and $Q_{2}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q_{2}})$ and a policy $\pi(\bm{y}_{k};\bm{\theta}_{\pi})$ by means of DNNs. However, the parameters of the main networks are now learned using three hypernetworks $h_{Q_{1}}(\bm{z}_{k};\bm{\theta}_{h_{Q_{1}}})$, $h_{Q_{2}}(\bm{z}_{k};\bm{\theta}_{h_{Q_{2}}})$, and $h_{\pi}(\bm{z}_{k};\bm{\theta}_{h_{\pi}})$:
$\begin{split}\bm{\theta}_{Q_{1}}&=h_{Q_{1}}(\bm{z}_{k};\bm{\theta}_{h_{Q_{1}}})\,,\\ \bm{\theta}_{Q_{2}}&=h_{Q_{2}}(\bm{z}_{k};\bm{\theta}_{h_{Q_{2}}})\,,\\ \bm{\theta}_{\pi}&=h_{\pi}(\bm{z}_{k};\bm{\theta}_{h_{\pi}})\,,\\ \end{split}$ (26)
where $\bm{z}_{k}$ indicates the hypernetwork input. With reference to Figure 2, we propose two variations of HypeRL: (i) we utilize the PDE parameter vector $\bm{\mu}$ as input to the hypernetworks, i.e., $\bm{z}_{k}=\bm{\mu}$ and (ii) we utilize state and PDE parameters as hypernetwork input, i.e., $\bm{z}_{k}=[\bm{y}_{k},\bm{\mu}]$.
Figure 2: Hyper policy and hyper value function architectures. In Figure 2(a) and 2(c) only the PDE parameters are used as input to the hypernetwork, i.e., $\bm{z}_{k}=\bm{\mu}$, while in Figure 2(b) and 2(d) the PDE state and parameters are used as input to the hypernetwork, i.e., $\bm{z}_{k}=[\bm{y}_{k},\bm{\mu}]$.
In this way, we are able to learn a representation of PDE parameters or of states and PDE parameter vector and encode this information in the parameters of the main networks. We can then obtain the actions from the policy as:
$\begin{split}\bm{\theta}_{\pi}&=h_{\pi}(\bm{z}_{k};\bm{\theta}_{h_{\pi}})\,,\\ \bm{u}_{k}&=\pi(\bm{y}_{k};\bm{\theta}_{\pi})\,,\\ \end{split}$ (27)
and analogously, we can obtain the Q-values as:
$\begin{split}\bm{\theta}_{Q_{i}}&=h_{Q_{i}}(\bm{z}_{k};\bm{\theta}_{h_{Q_{i}}})\,,\\ q_{i,t}&=Q_{i}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q_{i}})\,,\\ \end{split}$ (28)
where $q_{i,t}$ is the predicted Q-value by the action-value function $Q_{i}(\cdot,\cdot)$.
@@
The hypernetwork parameters are jointly optimized with the main network parameters by simply allowing the gradient of the TD3 training objectives [^5] We do not need to change the TD3 working principles and losses. (see in Equation (17) and (19)) to flow through the hypernetworks:
$\begin{split}\mathcal{L}(\bm{\theta}_{Q_{i}},\bm{\theta}_{h_{Q_{i}}})&=\mathbb{E}_{\bm{y}_{k},\bm{u}_{k},r_{k},\bm{y}_{k+1}\sim\mathcal{D}}[(r_{k}+\gamma\min_{i=1,2}\bar{Q}_{i}(\bm{y}_{k+1},\bm{u}_{k+1};\bm{\theta}_{\bar{Q}_{i}})-Q_{i}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q_{i}}))^{2}]\,,\\ \mathcal{L}(\bm{\theta}_{\pi},\bm{\theta}_{h_{\pi}})&=\mathbb{E}_{\bm{y}_{k}\sim\mathcal{D}}[-\nabla_{\bm{u}_{k}}Q_{1}(\bm{y}_{k},\pi(\bm{y}_{k};\bm{\theta}_{\pi});\bm{\theta}_{Q_{1}})].\end{split}$ (29)
@@
In Algorithm 2, we show the HypeRL TD3 pseudo code, where we highlight in blue the changes to the original TD3 algorithm. It is worth highlighting that our method can be plugged in any RL algorithm as we did not change their working principles.
Algorithm 2 Parameter-Informed HypeRL TD3
Initialize $Q_{1}(\bm{y},\bm{u};\bm{\theta}_{Q_{1}})$, $Q_{2}(\bm{y},\bm{u};\bm{\theta}_{Q_{2}})$, and $\pi(\bm{y};\bm{\theta}_{\pi})$
Initialize hypernetworks $h_{Q_{1}}(\bm{z};\bm{\theta}_{h_{Q_{1}}})$, $h_{Q_{2}}(\bm{z};\bm{\theta}_{h_{Q_{2}}})$, and $h_{\pi}(\bm{z};\bm{\theta}_{h_{\pi}})$ with random parameters $\bm{\theta}_{h_{Q_{1}}},\bm{\theta}_{h_{Q_{2}}},\bm{\theta}_{h_{\pi}}$
@@
Initialize memory buffer $\mathcal{D}$
@@
         Sample action $\bm{u}_{k}\sim\pi(\bm{y}_{k};\bm{\theta}_{\pi})+\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{\sigma})$
         Observe reward $r_{k}$ and new state $\bm{y}_{k+1}$
         Store tuple $(\bm{y}_{k},\bm{u}_{k},r_{k},\bm{y}_{k+1},\bm{\mu})$ in $\mathcal{D}$
@@
              Sample mini-batch $(\bm{y}_{k},\bm{u},r,\bm{y}_{k+1},\bm{\mu})$ from $\mathcal{D}$
               Set $\bm{z}_{k+1}=\bm{\mu}$ or $\bm{z}_{k+1}=[\bm{y}_{k+1},\bm{\mu}]$
@@
              $\bm{u}_{k+1}\leftarrow\bar{\pi}(\bm{y}_{k+1};\bm{\theta}_{\bar{\pi}})+\bm{\epsilon}$, where $\bm{\epsilon}\sim\text{clip}(\mathcal{N}(\bm{0},\bar{\bm{\sigma}}),-c,c)$
@@
              $q^{k}\leftarrow r_{k}+\gamma\min_{i=1,2}\bar{Q}(\bm{y}_{k+1},\bm{u}_{k+1};\bm{\theta}_{\bar{Q}_{i}})$
@@
              $\mathcal{L}(\bm{\theta}_{Q_{i}},{\color[rgb]{0,0,1}\bm{\theta}_{h_{Q_{i}}}})=\mathbb{E}_{\bm{y}_{k},\bm{u}_{k},r_{k},\bm{y}_{k+1}\sim\mathcal{D}}[(q_{k}-Q_{i}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q_{i}}))^{2}]$     with    $i\in\{1,2\}$
@@
                  $\mathcal{L}(\bm{\theta}_{\pi},{\color[rgb]{0,0,1}\bm{\theta}_{h_{\pi}}})=\mathbb{E}_{\bm{y}_{k}\sim\mathcal{D}}[-\nabla_{\bm{u}_{k}}Q_{1}(\bm{y}_{k},\pi(\bm{y}_{k};\bm{\theta}_{\pi});\bm{\theta}_{Q_{1}})$

### v2

@@ 2.2.1 Twin-Delayed Deep Deterministic Policy Gradient
[…]
@@
$Q(\bm{y}_{k},\bm{u}_{k})\leftarrow Q(\bm{y}_{k},\bm{u}_{k})+\alpha\Big(r_{k+1}+\gamma\max_{\bm{u}_{k}\in U_{\text{ad}}}Q(\bm{y}_{k+1},\cdot)-Q(\bm{y}_{k},\bm{u}_{k})\Big)\,$, (11)
@@
$Q(\bm{y}_{k},\bm{u}_{k})\leftarrow Q(\bm{y}_{k},\bm{u}_{k})+\alpha\Big(r_{k+1}+\gamma Q(\bm{y}_{k+1},\bm{u}_{k+1})-Q(\bm{y}_{k},\bm{u}_{k})\Big)\,$, (12)
@@
TD3 relies on a memory buffer $\mathcal{M}$ to store the interaction data $(\bm{y}_{k},\bm{u}_{k},r_{k+1},\bm{y}_{k+1})$ for all time-steps $k$. Given a randomly-sampled batch of interaction tuples, we can employ Equation (12) as a loss function for updating the parameters $\bm{\theta}_{Q}$ of the action-value function $Q(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q})$ as:
[…]
where the so-called target networks $\bar{Q}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{\bar{Q}})$ and $\bar{\pi}(\bm{y}_{k};\bm{\theta}_{\bar{\pi}})$ are copies of $Q(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q})$ and $\pi(\bm{y}_{k};\bm{\theta}_{\pi})$, respectively, with frozen parameters, i.e., they are not updated in the backpropagation step to improve the stability of the training. We indicate with $\bm{\epsilon}\sim\text{clip}(\mathcal{N}(\bm{0},\bar{\bm{\sigma}}),-c,c)$ the noise added to estimate the action value in the interval [c,c][-c,c] around the target action. To reduce the problem of overestimation of the target Q-values, TD3 estimates two independent action-value functions, namely $Q_{1}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q_{1}})$ and $Q_{2}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{Q_{2}})$, and two target action-value functions $\bar{Q}_{1}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{\bar{Q}_{1}})$ and $\bar{Q}_{2}(\bm{y}_{k},\bm{u}_{k};\bm{\theta}_{\bar{Q}_{2}})$, and computes the target value for regression (see (13)) as:
$\underbrace{r_{k+1}+\gamma\min_{i=1,2}\bar{Q}_{i}(\bm{y}_{k+1},\bm{u}_{k+1};\bm{\theta}_{\bar{Q}_{i}}).}_{\text{target value}}$
@@
$\mathcal{L}(\bm{\theta}_{\pi})=\mathbb{E}_{\bm{y}_{k}\sim\mathcal{M}}[-\nabla_{\bm{u}_{k}}Q_{1}(\bm{y}_{k},\pi(\bm{y}_{k};\bm{\theta}_{\pi});\bm{\theta}_{Q_{1}})].$
@@
[…]
@@ 2.3 Hypernetworks
A hypernetwork [22] is a neural network (NN) that generates the weights and biases of another NN, often referred to as main or primary network. Formally, a hypernetwork $h:Z\subset\mathbb{R}^{|\bm{z}|}\to\mathbb{R}^{|\bm{\theta}_{f}|}$ learns the parameters $\bm{\theta}_{f}$ of the main network $f:X\subset\mathbb{R}^{|\bm{x}|}\to W\subset\mathbb{R}^{|\bm{w}|}$. If we consider a standard supervised learning regression task, and we assume the availability of a dataset of $N$ input-output pairs $(\bm{x}^{(1)},\bm{w}^{(1)}),\cdots,(\bm{x}^{(N)},\bm{w}^{(N)})]$ and hypernetwork inputs $\bm{z}^{(1)},\ldots,\bm{z}^{(N)}]$, we can write:
[…]
@@ 3 Methodology for HypeRL
Given all the elements introduced in the previous section, we are now ready to set our proposed strategy to address OC of parametric dynamical systems through RL.
@@ 3.1 Problem Settings
We cast the OC problem of parametric dynamical systems, introduced in Equation (1), as RL problem, where an agent, i.e., the controller, aims to learn the OC strategy by interacting with an unknown environment, governed by the dynamical system. Differently from the majority of the literature on RL for OC of dynamical systems, we focus on devising a control strategy that is adaptable to changes of the systems’ dynamics deriving from variations of known parameters $\bm{\mu}$. The RL environment is defined by a transition and reward functions:
@@
where the transition and reward function are assumed to be deterministic and dependent on the parameter vector $\bm{\mu}$. The parameter vector $\bm{\mu}$ might describe a physical parameter of the system, a reference value, or a navigation target. The transition function corresponds to the fully-discretized form of the state equation (see Equation (1)) obtained by introducing a suitable time-integration scheme over a partition of $t_{0},t_{f}]$ made by $N_{t}$ time-steps $\{t_{k}\}_{k=0}^{N_{t}}$ such that the step size is $\Delta t=(t_{f}-t_{0})/N_{t}$. For example, by using an explicit Runge-Kutta scheme, we obtain the following full-order model:
$\begin{split}\bm{y}_{k+1}=\underbrace{\bm{y}_{k}+\Delta t\Phi(t_{k},\bm{y}_{k},\bm{u}_{k};\Delta t,\bm{F},\bm{\mu})}_{:=\,T(\bm{y}_{k},\bm{u}_{k};\bm{\mu})},\qquad k=0,\ldots,N_{t}-1,\end{split}$
where $\bm{y}_{k}\approx\bm{y}(t_{k})$ and $\bm{u}_{k}\approx\bm{u}(t_{k})$, and $\Phi$ denotes the integration method’s increment function related to the state equation. The agent may observe the full state $\bm{y}_{k}$ or only a part of it. In addition, we assume accurate knowledge of the parameter $\bm{\mu}$. While the knowledge of physical parameters might seem a strong hypothesis, the advances of machine learning have opened new doors toward estimating such parameters from data with high accuracy, e.g., [55]. We indicate the agent’s state at timestep $k$ with $\bm{z}_{k}=[\bm{y}_{k},\bm{\mu}]$.
With reference to Algorithm 1, the agent-environment interaction scheme is organized in two loops: the outer loop indicates the training episode; instead, the inner loop refers to the number of time-steps for each episode. At the beginning of each new episode we sample a random initial condition and a PDE parameters value $\bm{\mu}$ to obtain the initial state. The agent utilizes the state at the current time-step and the PDE parameters to select a control input. The control input is fed to the transition and reward models to obtain the state at the next time-step and the reward. The tuples $(\bm{z}_{k},\bm{u}_{k},r_{k},\bm{z}_{k+1})$ collected at each time-step $k$ of the interaction are used to train the policy (and value function) deep NN to maximize the expected cumulative reward at each episode.
Algorithm 1 Episodic RL for control of parametric dynamical system
Initialize policy and value function parameters $\bm{\theta}_{\pi}$ and $\bm{\theta}_{Q}$
@@
  Sample an initial condition and input parameter $\bm{\mu}$
@@
   Sample action from a policy $\bm{u}_{k}\sim\pi(\bm{z}_{k};\bm{\theta}_{\pi})+\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{\sigma})$
   Observe reward $r_{k+1}=R(\bm{y}_{k},\bm{u}_{k};\bm{\mu})$ and new state $\bm{y}_{k+1}=T(\bm{y}_{k},\bm{u}_{k};\bm{\mu})$
@@
     Update policy and value function using the tuple $(\bm{z}_{k},\bm{u}_{k},r_{k+1},\bm{z}_{k+1})$
@@
In this setting, we present a novel RL framework that can efficiently learn control policies for parametric PDEs from limited samples and that can generalize to new, unseen instances of the systems’ parameters $\bm{\mu}$ (see Figure 1). In contrast with the widely-used concatenation of information in the agent’s state, to enhance the sample-efficiency and generalization capabilities of RL agents, we propose a parameter-informed HypeRL architecture relying on hypernetworks (see Section 2.3). Hypernetworks allow us to express the weights and biases of the value function and policy as functions of $\bm{y}_{k}$ and $\bm{\mu}$. This new paradigm for encoding the information of physical and task-dependent parameters drastically improves the performance of the RL agent in terms of total cumulative reward, sample efficiency, and generalization with respect to state-of-the-art RL approaches.
3.2 HypeRL-TD3
@@
In this work, we enhance the TD3 algorithm (see Section 2.2.1) with hypernetworks. However, our method can be easily and directly applied to other RL algorithms, such as PPO and SAC. Analogously to the TD3 algorithm, we rely on the estimation of two action-value functions $Q_{1}(\bm{z}_{k},\bm{u}_{k};\bm{\theta}_{Q_{1}})$ and $Q_{2}(\bm{z}_{k},\bm{u}_{k};\bm{\theta}_{Q_{2}})$ and a policy $\pi(\bm{z}_{k};\bm{\theta}_{\pi})$ by means of deep NNs. However, the parameters of the main networks are now learned using three hypernetworks $h_{Q_{1}}(\bm{z}_{k};\bm{\theta}_{h_{Q_{1}}})$, $h_{Q_{2}}(\bm{z}_{k};\bm{\theta}_{h_{Q_{2}}})$, and $h_{\pi}(\bm{z}_{k};\bm{\theta}_{h_{\pi}})$. In this way, we are able to learn a representation of state $\bm{y}_{k}$ and input parameter vector $\bm{\mu}$ and encode this information in the weights and biases of the main networks. The hyper-policy is defined as:
[…]
and analogously, the hyper-value functions as:
[…]
@@
where $q_{i,k}$ is the predicted Q-value by the action-value function $Q_{i}(\cdot,\cdot)$. The hypernetwork parameters are jointly optimized with the main network parameters by simply allowing the gradient of the loss functions to flow through the hypernetworks. Policy and value functions are updated using the TD3 training objectives:
[…]
where, similarly to [18], we found beneficial to the agent’s performance replacing the mean-squared error with the Huber loss to update the hyper-value function.
@@
In Algorithm 2, we show the HypeRL-TD3 pseudo code, where we highlight in blue the changes to the original TD3 algorithm.
Algorithm 2 HypeRL-TD3
Initialize hypernetworks $h_{Q_{1}}(\cdot;\bm{\theta}_{h_{Q_{1}}})$, $h_{Q_{2}}(\cdot;\bm{\theta}_{h_{Q_{2}}})$, and $h_{\pi}(\bm{z};\bm{\theta}_{h_{\pi}})$ with random parameters $\bm{\theta}_{h_{Q_{1}}},\bm{\theta}_{h_{Q_{2}}},\bm{\theta}_{h_{\pi}}$
@@
Initialize main networks $Q_{1}(\cdot,\cdot;\bm{\theta}_{Q_{1}})$, $Q_{2}(\cdot,\cdot;\bm{\theta}_{Q_{2}})$, and $\pi(\cdot;\bm{\theta}_{\pi})$
Initialize target networks $\bar{Q}_{1}(\cdot,\cdot;\bm{\theta}_{\bar{Q}_{1}})$, $\bar{Q}_{2}(\cdot,\cdot;\bm{\theta}_{\bar{Q}_{2}})$, and $\bar{\pi}(\cdot;\bm{\theta}_{\bar{\pi}})$
Initialize memory buffer $\mathcal{M}$
@@
   Sample action $\bm{u}_{k}\sim\pi(\bm{z}_{k};\bm{\theta}_{\pi})+\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{\sigma})$
   Observe reward $r_{k+1}$ and new state $\bm{y}_{k+1}$
   Store tuple $(\bm{z}_{k},\bm{u}_{k},r_{k+1},\bm{z}_{k+1})$ in $\mathcal{M}$
@@
     Sample mini-batch $(\bm{z}_{k},\bm{u},r_{k+1},\bm{z}_{k+1})$ from $\mathcal{M}$
@@
     $\bm{u}_{k+1}\leftarrow\bar{\pi}(\bm{z}_{k+1};\bm{\theta}_{\bar{\pi}})+\bm{\epsilon}$, where $\bm{\epsilon}\sim\text{clip}(\mathcal{N}(\bm{0},\bar{\bm{\sigma}}),-c,c)$
@@
     $q^{k}\leftarrow r_{k+1}+\gamma\min_{i=1,2}\bar{Q}(\bm{z}_{k+1},\bm{u}_{k+1};\bm{\theta}_{\bar{Q}_{i}})$
@@
     $\mathcal{L}(\bm{\theta}_{Q_{i}},{\color[rgb]{0,0,1}\bm{\theta}_{h_{Q_{i}}}})=\mathbb{E}_{\bm{z}_{k},\bm{u}_{k},r_{k+1},\bm{z}_{k+1}\sim\mathcal{M}}[{\color[rgb]{0,0,1}\texttt{Huber}}(q_{k}-Q_{i}(\bm{z}_{k},\bm{u}_{k};\bm{\theta}_{Q_{i}}))]$ with $i\in\{1,2\}$
@@
      $\mathcal{L}(\bm{\theta}_{\pi},{\color[rgb]{0,0,1}\bm{\theta}_{h_{\pi}}})=\mathbb{E}_{\bm{z}_{k}\sim\mathcal{M}}[-\nabla_{\bm{u}_{k}}Q_{1}(\bm{z}_{k},\pi(\bm{z}_{k};\bm{\theta}_{\pi});\bm{\theta}_{Q_{1}})]$